# Bridging Semantic Understanding and Popularity Bias with LLMs

### v2

@@ 1. Introduction
As recommender systems strive to become more trustworthy, enhancing semantic understanding of core echical concepts, such as popularity bias, is essential for advancing cognitive capabilities of web applications (Zhang et al., 2025d; Gao et al., 2025). Popularity bias occurs when popular items are recommended too frequently, while other products that may deserve more attention are ignored (Abdollahpouri et al., 2019). This surface-level issue, however, masks a more complex interaction between user preferences, item rankings, and recommendation algorithms (Tang et al., 2024). In this context, semantic understanding refers to a model’s ability to grasp the deeper implications of popularity bias, rather than merely recognizing its surface-level features (Xu, 2025). In other words, a comprehensive semantic understanding of popularity bias requires recognizing not only the observable imbalance in recommendations but also the underlying causal mechanisms that give rise to it.
@@
Developing a deeper semantic understanding of popularity bias is key to addressing challenges such as the long-tail problem (Zhang et al., 2025c), product homogeneity (Lin et al., 2025), and personalized user satisfaction (Silva et al., 2021; Zhang et al., 2025a). Existing methods, such as PAAC (Cai et al., 2024), primarily focus on surface-level mitigation by increasing recommendation diversity or boosting exposure to non-popular items. Although terms like “diversity” and “debiasing” are frequently used in prompts, our investigation with Large Language Models (LLMs) reveals that such cues remain shallow representations of popularity bias (Klimashevskaia et al., 2024; Xiang et al., 2025). They fail to encode the causal semantics that link user behavior, item features, and system design to the emergence of bias.
@@ 2.1. Semantic Understanding for Recommender Systems
As web applications continue to evolve, achieving a deeper semantic understanding of complex definitions is crucial for enhancing the cognitive capabilities of models (Zhang et al., 2024b, 2025b). Bhattacharya and Pandey (Bhattacharya and Pandey, 2024) developed an agricultural ontology for extracting relationships from text using Natural Language Processing, aiming to improve semantic understanding. Wang et al. (Wang, 2024) proposed a deep learning model based on the Transformer architecture, integrating a self-attention mechanism to capture long-distance dependencies in text for improved semantic understanding tasks.
Several researchers have focused specifically on semantic understanding for recommender systems (Liu et al., 2025; Zhang et al., 2025e). Shang et al. (Shang et al., 2024) used LLMs to integrate semantic understanding with user preferences, combining a fine-tuned Roberta semantic analysis model with a multi-modal user preference extraction mechanism. Similarly, Wan et al. (Wan et al., 2024) leveraged LLMs for semantic understanding by incorporating real-time scene information into recommender systems, enhancing RecLLM efficiency without requiring the model to process entire real-time scene texts directly.
Despite these advancements, a significant gap remains in the current understanding of semantics: the lack of explicit attention to ethical definitions, such as popularity bias (Zhang et al., 2024a). While LLMs excel at generating semantically relevant recommendations, the training data often reflects existing popularity distributions, leading to the overexposure of popular items and the neglect of niche content (Nie et al., 2024). This imbalance can result in unfair outcomes for users with diverse preferences, negatively impacting both fairness and user satisfaction. Recent studies have started exploring fairness in RecLLM through fine-tuning and related techniques (Jiang et al., 2024; Hua et al., 2024), primarily focusing on group fairness (e.g., ensuring parity across demographic groups). However, the semantic understanding of popularity bias, particularly its impact on the interaction between items and users, remains largely unaddressed, leaving a significant gap in the current research landscape.
@@ 2.2. Popularity Bias
Research has focused on mitigating popularity bias in traditional recommender systems (Chen et al., 2024; Ungruh et al., 2024). Some approaches explicitly diversify recommendation lists by promoting long-tail items, while others modify learning objectives to penalize the recommendation of popular items (Huang et al., 2024). These methods have shown varying levels of success in mitigating bias in traditional settings. However, the direct application of these conventional debiasing techniques to the RecLLM paradigm is not straightforward due to the distinct architectural and operational characteristics of LLMs. LLMs’ black-box nature and their reliance on vast amounts of pre-trained data make it difficult to apply traditional debiasing algorithms directly (Dai et al., 2025). Moreover, in the emerging field of RecLLM, the study of popularity bias and its mitigation remains underdeveloped. Existing attempts to address this issue in LLMs often rely on basic data partitioning or simplistic re-ranking strategies based on item popularity, which fall short of comprehensively tackling the problem (Ortega et al., 2024; Luo et al., 2025).
@@ 3.2. Item and User Popularity Categorization
To inform the debiasing methodology, we establish a systematic categorization of both items and users based on popularity. This process begins with item classification. Leveraging the Pareto principle (i.e., the $20/80$ rule), items are sorted by their interaction counts (Bender, 1981). Empirical evidence from long-tail distributions shows that a small fraction of items typically accounts for the majority of interactions. Accordingly, the top $20\%$ of items with the highest interaction counts are designated as P (popular items or hot items), while the remaining $80\%$ are classified as N (niche items or tail items). This $2/8$ division is consistent with the statistical characteristics of real-world recommender data, where item consumption naturally follows a long-tail distribution (Zhao et al., 2025). The distinction between popular and niche items thus provides a principled basis for analyzing item popularity and guiding debiasing strategies, enabling the model to differentiate between widely consumed and less frequently accessed content.

### v3

@@ 1. Introduction
As recommender systems strive to become more trustworthy, enhancing semantic understanding of core echical concepts, such as popularity bias, is essential for advancing cognitive capabilities of web applications (Zhang et al., 2025c; Gao et al., 2025). Popularity bias occurs when popular items are recommended too frequently, while other products that may deserve more attention are ignored (Abdollahpouri et al., 2019). This surface-level issue, however, masks a more complex interaction between user preferences, item rankings, and recommendation algorithms (Tang et al., 2024; Zhang et al., 2025a). In this context, semantic understanding refers to a model’s ability to grasp the deeper implications of popularity bias, rather than merely recognizing its surface-level features (Xu, 2025). In other words, a comprehensive semantic understanding of popularity bias requires recognizing not only the observable imbalance in recommendations but also the underlying causal mechanisms that give rise to it.
@@
Developing a deeper semantic understanding of popularity bias is key to addressing challenges such as the long-tail problem (Zhang et al., 2025b), product homogeneity (Lin et al., 2025), and personalized user satisfaction (Silva et al., 2021; Zhang et al., 2025d). Existing methods, such as PAAC (Cai et al., 2024), primarily focus on surface-level mitigation by increasing recommendation diversity or boosting exposure to non-popular items. Although terms like “diversity” and “debiasing” are frequently used in prompts, our investigation with Large Language Models (LLMs) reveals that such cues remain shallow representations of popularity bias (Klimashevskaia et al., 2024; Xiang et al., 2025). They fail to encode the causal semantics that link user behavior, item features, and system design to the emergence of bias.
@@ 2.1. Semantic Understanding for Recommender Systems
As web applications continue to evolve, achieving a deeper semantic understanding of complex definitions is crucial for enhancing the cognitive capabilities of models (Zhang et al., 2024a, 2025e; Yang et al., 2025). Bhattacharya and Pandey (Bhattacharya and Pandey, 2024) developed an agricultural ontology for extracting relationships from text using Natural Language Processing, aiming to improve semantic understanding. Wang et al. (Wang, 2024) proposed a deep learning model based on the Transformer architecture, integrating a self-attention mechanism to capture long-distance dependencies in text for improved semantic understanding tasks.
Several researchers have focused specifically on semantic understanding for recommender systems (Liu et al., 2025; Zhang et al., 2025f; Wanyan et al., 2025; Ma et al., 2025). Shang et al. (Shang et al., 2024) used LLMs to integrate semantic understanding with user preferences, combining a fine-tuned Roberta semantic analysis model with a multi-modal user preference extraction mechanism. Similarly, Wan et al. (Wan et al., 2024) leveraged LLMs for semantic understanding by incorporating real-time scene information into recommender systems, enhancing RecLLM efficiency without requiring the model to process entire real-time scene texts directly.
Despite these advancements, a significant gap remains in the current understanding of semantics: the lack of explicit attention to ethical definitions, such as popularity bias (Zhang et al., 2024b). While LLMs excel at generating semantically relevant recommendations, the training data often reflects existing popularity distributions, leading to the overexposure of popular items and the neglect of niche content (Nie et al., 2024). This imbalance can result in unfair outcomes for users with diverse preferences, negatively impacting both fairness and user satisfaction. Recent studies have started exploring fairness in RecLLM through fine-tuning and related techniques (Jiang et al., 2024; Hua et al., 2024), primarily focusing on group fairness (e.g., ensuring parity across demographic groups). However, the semantic understanding of popularity bias, particularly its impact on the interaction between items and users, remains largely unaddressed, leaving a significant gap in the current research landscape.
@@ 2.2. Popularity Bias
Research has focused on mitigating popularity bias in traditional recommender systems (Chen et al., 2024; Ungruh et al., 2024). Some approaches explicitly diversify recommendation lists by promoting long-tail items, while others modify learning objectives to penalize the recommendation of popular items (Huang et al., 2024). These methods have shown varying levels of success in mitigating bias in traditional settings. However, the direct application of these conventional debiasing techniques to the RecLLM paradigm is not straightforward due to the distinct architectural and operational characteristics of LLMs. LLMs’ black-box nature and their reliance on vast amounts of pre-trained data make it difficult to apply traditional debiasing algorithms directly (Dai et al., 2025; Zhao et al., 2025a; Ren et al., 2025). Moreover, in the emerging field of RecLLM, the study of popularity bias and its mitigation remains underdeveloped. Existing attempts to address this issue in LLMs often rely on basic data partitioning or simplistic re-ranking strategies based on item popularity, which fall short of comprehensively tackling the problem (Ortega et al., 2024; Luo et al., 2025).
@@ 3.2. Item and User Popularity Categorization
To inform the debiasing methodology, we establish a systematic categorization of both items and users based on popularity. This process begins with item classification. Leveraging the Pareto principle (i.e., the $20/80$ rule), items are sorted by their interaction counts (Bender, 1981). Empirical evidence from long-tail distributions shows that a small fraction of items typically accounts for the majority of interactions. Accordingly, the top $20\%$ of items with the highest interaction counts are designated as P (popular items or hot items), while the remaining $80\%$ are classified as N (niche items or tail items). This $2/8$ division is consistent with the statistical characteristics of real-world recommender data, where item consumption naturally follows a long-tail distribution (Zhao et al., 2025b). The distinction between popular and niche items thus provides a principled basis for analyzing item popularity and guiding debiasing strategies, enabling the model to differentiate between widely consumed and less frequently accessed content.